OpenStudy (anonymous):

what would the drummer do to make the sound of the drum give a note of lower pitch

OpenStudy (chenping):

pitch = frequency. If it's an electrical drum, lower the power supply :P If not, hit softer >< kidding.. juz adjust the instrument, nothing else.. ==

OpenStudy (shane_b):

He'd have to loosen the drum skin which would reduce the frequency of the oscillation when he hits the drum. Basically the same concept as loosening a guitar string.

OpenStudy (anonymous):

for a string \[f = 1/2l \times \sqrt{T/u}\] f- frequency l-length T- tension u-mass per unit length Similarly even for a drum, you can reduce tension i.e loosen the drum skin..

OpenStudy (anonymous):

loosening the drum skin will lower the viberating frequency nd hence produce low pitch note